# Classifying localizing subcategories of a Grothendieck category

### v1

@@ 2. Categories with enough atoms
If $\sigma$ is a limit ordinal, then $\mathcal{A}_{\sigma}$ is generated by $\bigcup_{\rho<\sigma}\mathcal{A}_{\rho}$ and so we may assume that $\sigma$ is a non-limit odinal. Let $\mathcal{C}$ be the subclass of all $\delta$-critical objects in $\mathcal{A}$ with $\delta\leq\sigma$ and we have to show that $\mathcal{A}_{\sigma}=\left\langle\mathcal{C}\right\rangle_{\rm loc}$. We prove the claim by transfinite induction on $\sigma$. The case $\sigma=0$ is clear and so we assume that $\sigma>0$. Let $\mathcal{D}$ be the class of $\sigma$-critical objects in $\mathcal{A}$. Then we have the following equalities
[…]
where the first equality holds by the definition and the second equality holds by [K, Proposition 4.18] and section 2 and the last equality holds using the induction hypothesis. It now follows from [K, Proposition 4.14] that $\mathcal{A}_{\sigma}=\left\langle\mathcal{C}\right\rangle_{\rm loc}$. ∎
[…]
If $M$ is a nonzero object in $\mathcal{A}$ with Krull-Gabriel dimension, then $M$ has a critical subobject (and so a monoform subobject)
@@
Since ordinals satisfy the descending chain condition, we can choose a nonzero subobject $N$ of $M$ of minimal Krull-Gabriel dimension $\sigma$. Clearly $\sigma$ is non-limit ordinal and $t_{\sigma-1}(N)=0$. Since $F_{\sigma-1}(N)\in(\mathcal{A}/\mathcal{A}_{\sigma-1})_{0}$, it follows from [St, Chap VI, Proposition 2.5] that $F_{\sigma-1}(N)$ contains a simple subobject $S$. Then $N$ contains a subobject $H$ such that $F_{\sigma-1}(H)=S$ by [Po, Chap 4, Corollary 3.10]. Now, section 2 implies that $H$ is a $\sigma$-critical. ∎
[…]
@@
If $M$ has Krull-Gabriel dimension, then every nonzero quotient of $M$ has Krull-Gabriel dimension as well. Thus section 2 implies that every quotient of $M$ contains a monoform subobject so that $M$ is locally monoform. The second assertion is clear. ∎
3. Locally finitely generated categories
For $\alpha\in\operatorname{ASpec}\mathcal{A}$ and monoform objects $H_{1}$ and $H_{2}$ in $\mathcal{A}$ satisfying $\alpha=\overline{H_{1}}=\overline{H_{2}}$, we have $E(H_{1})=E(H_{2})$. The isomorphism class of all such $E(H)$ such that $\alpha=\overline{H}$ is denoted by $E(\alpha)$. We observe that $E(\alpha)$ is an indecomposable injective object in $\mathcal{A}$. Because if $E(\alpha)=E(H)$ for some monoform object $H$ in $\mathcal{A}$ and $E(\alpha)=E_{1}\oplus E_{2}$, then $E_{1}\cap H$ and $E_{2}\cap H$ are nonzero monoform subobjecs of $H$. Thus $E_{1}\cap H\cap E_{2}\cap H$ is nonzero which is a contradiction. We show that for any object $M$ in $\mathcal{A}$, $\operatorname{ASupp}M$ can be determined in terms of indecomposable injective objects.
[…]
@@ 3. Locally finitely generated categories
Let $\mathcal{X}$ be a localizing subcategory of $\mathcal{A}$ with the canonical functor $F:\mathcal{A}\rightarrow\mathcal{A}/\mathcal{X}$ and the right adjoint $G:\mathcal{A}\rightarrow\mathcal{A}/\mathcal{X}$. We recall from [GL] that the right perpendicular category $\mathcal{X}^{\bot}$ of $\mathcal{X}$ is the full subcategory of all objects $M\in\mathcal{A}$ satisfying $\operatorname{Hom}(X,M)=\operatorname{Ext}^{1}(X,M)=0$ for all $X\in\mathcal{X}$. For each object $M$, the adjointness provides a natural morphism $\eta_{M}:M\rightarrow GF(M)$ such that $t_{\mathcal{X}}(M)=\operatorname{Ker}\eta_{M},\operatorname{Cok}\eta_{M}\in\mathcal{X}$. The morphism $\eta_{M}$ is said to be the localization of $M$ in $\mathcal{X}$ and the functor $F(-)$ is denoted by $(-)_{\mathcal{X}}$. It is known that an object $M$ is in $\mathcal{X}^{\bot}$ if and only if $\eta_{M}$ is an isommorphism. Furthermore, according to [GL, Proposition 2.2], the functor $F:\mathcal{X}^{\bot}\rightarrow\mathcal{A}/\mathcal{X}$ is a natural equivalence. Then we can consider $\mathcal{A}/\mathcal{X}$ as a full subcategory $\mathcal{A}$. Again, by virtue of [GL, Proposition 2.2], the functor $GF(-)=G((-)_{\mathcal{X}}):\mathcal{A}\rightarrow\mathcal{X}^{\bot}$ is a left adjoint of the inclusion $\mathcal{X}^{\bot}\subseteq\mathcal{A}$. This implies that the localization functor preserves colimits.
The collection of the isomorphism classes of indecomposable injective objects in $\mathcal{A}$, denoted by $\operatorname{Sp}\mathcal{A}$, forms a set. This is because every indecomposable injective object is the injective envelope of some quotient of an element of the generating set of $\mathcal{A}$.
[…]
(1) For any indecomposable injective object $I$ in $\operatorname{Sp}\mathcal{A}$, the localizing subcategory $\mathcal{X}(I)=\{M\in\mathcal{A}|\hskip 2.84544pt\operatorname{Hom}(M,I)=0\}$ admits a canonical exact functor $(-)_{I}=(-)_{\mathcal{X}(I)}:\mathcal{A}\rightarrow\mathcal{A}/\mathcal{X}(I)$. The image of every object $M$ under this functor is said to be the localization of $M$ at $I$ and we denote it by $M_{I}$.
(2) The Ziegler support $\operatorname{ZSupp}(M)$ of an object $M$ in $\mathcal{A}$ is a subset of $\operatorname{Sp}\mathcal{A}$, that is
$\operatorname{ZSupp}M=\{I\in\operatorname{Zg}\mathcal{A}|\hskip 2.84544ptM_{I}\neq 0\}.$
The definition forces that $\operatorname{ZSupp}M=\{I\in\operatorname{Zg}\mathcal{A}|\hskip 2.84544pt\operatorname{Hom}(M,I)\neq 0\}$. Since $(-)_{I}$ is an exact functor, for any exact sequence $0\rightarrow N\rightarrow M\rightarrow K\rightarrow 0$ of objects in $\mathcal{A}$, we have
$\operatorname{ZSupp}M=\operatorname{ZSupp}N\cup\operatorname{ZSupp}K.$
[…]
Let $\{M_{\lambda}|\hskip 2.84544pt\lambda\in\varLambda\}$ be a family of objects in $\mathcal{A}$. Then $\operatorname{ZSupp}(\bigoplus_{\lambda\in\varLambda}M_{\lambda})=\bigcup_{\lambda\in\varLambda}\operatorname{ZSupp}M_{\lambda}$. In particular, $\operatorname{ZSupp}^{-1}\mathcal{U}$ is a localizing subcategory of $\mathcal{A}$ for any subset $\mathcal{U}$ of $\operatorname{Sp}\mathcal{A}$.
Proof.
The inclusion $\bigcup_{\lambda\in\varLambda}\operatorname{ZSupp}M_{\lambda}\subseteq\operatorname{ZSupp}(\bigoplus_{\lambda\in\varLambda}M_{\lambda})$ is clear. For $I\in\operatorname{ZSupp}(\bigoplus_{\lambda\in\varLambda}M_{\lambda})$, the isomorphism $\operatorname{Hom}(\bigoplus_{\lambda\in\varLambda}M_{\lambda},I)\cong\prod_{\lambda\in\varLambda}\operatorname{Hom}(M_{\lambda},I)$ implies that $I\in\bigcup_{\lambda\in\varLambda}\operatorname{ZSupp}M_{\lambda}$. The second assertion follows from the first part. ∎
[…]
The set $\{\operatorname{ZSupp}M|\hskip 2.84544ptM\in\mathcal{A}\}$ forms a basis of open subsets of $\operatorname{Sp}\mathcal{A}$ for a topology on $\operatorname{Sp}\mathcal{A}$.
Proof.
For any indecomposable injective object $I\in\operatorname{Sp}\mathcal{A}$, we have $I\in\operatorname{ZSupp}I$. Moreover, for any objects $M_{1}$ and $M_{2}$ in $\mathcal{A}$ and $I\in\operatorname{ZSupp}M_{1}\cap\operatorname{ZSupp}M_{2}$, there exist nonzero homomorphisms $f_{i}:M_{i}\rightarrow I$ for $i=1,2$. Since $I$ is a uniform object in $\mathcal{A}$, the object $N=\operatorname{Im}f_{1}\cap\operatorname{Im}f_{2}$ is nonzero and we have $I\in\operatorname{ZSupp}N\subseteq\operatorname{ZSupp}M_{1}\cap\operatorname{ZSupp}M_{2}$. ∎
For any subcategory $\mathcal{X}$ of $\mathcal{A}$, we define $\operatorname{ZSupp}\mathcal{X}=\bigcup_{M\in\mathcal{X}}\operatorname{ZSupp}M$. Moreover, for any subset $\mathcal{U}$ of $\operatorname{Zg}\mathcal{A}$, we define $\operatorname{ZSupp}^{-1}\mathcal{U}=\{M\in\mathcal{A}|\hskip 2.84544pt\operatorname{ZSupp}M\subset\mathcal{U}\}$. It is clear that $\operatorname{ZSupp}^{-1}\mathcal{U}$ is a Serre subcategory of $\mathcal{A}$.
[…]
The map $\theta:\operatorname{ASpec}\mathcal{A}\rightarrow\operatorname{Sp}\mathcal{A}$, given by $\alpha\mapsto E(\alpha)$ embeds $\operatorname{ASpec}\mathcal{A}$ as a subspace of $\operatorname{Sp}\mathcal{A}$. In particular, if $\mathcal{A}$ is a locally monoform category, then $\theta$ is a homeomorphism.
Proof.
It is clear that $\theta$ is injective. Moreover, for any object $M$ in $\mathcal{A}$, it follows from section 3 that $\theta^{-1}(\operatorname{ZSupp}M)=\operatorname{ASupp}M$. To prove the second assertion, if $E$ is an indecomposable injective module, then it contains a monoform object $H$. Then $H=E(\alpha)$, where $\alpha=\overline{H}$. ∎
For any localizing subcategory $\mathcal{X}$ of $\mathcal{A}$, it turns out from [G] that $\mathcal{X}$ and $\mathcal{A}/\mathcal{X}$ are Grothendieck categories. The assignment $X\mapsto E(X)$ induces an injective map $E:\operatorname{Sp}\mathcal{X}\rightarrow\operatorname{Sp}\mathcal{A}$, where $E(X)$ is the injective envelope of $X$ in $\mathcal{A}$. Furthermore, $X\mapsto G(X)$ induces an injective map $G:\operatorname{Sp}(\mathcal{A}/\mathcal{X})\rightarrow\operatorname{Sp}\mathcal{A}$, where $G:\mathcal{A}/\mathcal{X}\rightarrow\mathcal{A}$ is the right adjoint of the canonical $F:\mathcal{A}\rightarrow\mathcal{A}/\mathcal{X}$. If we consider both maps as identification, then $\operatorname{Sp}\mathcal{A}=\operatorname{Sp}\mathcal{X}\cup\operatorname{Sp}(\mathcal{A}/\mathcal{X})$ and $\operatorname{Sp}\mathcal{X}\cap\operatorname{Sp}(\mathcal{A}/\mathcal{X})=\emptyset$ [G, III.3, Corollaire 2].
[…]
Let $\mathcal{X}$ be a localizing subcategory of $\mathcal{A}$. Then $\operatorname{Sp}\mathcal{X}$ is homeomorphic to the open subset $\operatorname{ZASupp}\mathcal{X}$ of $\operatorname{Sp}\mathcal{A}$.
Proof.
[…]
Let $\mathcal{X}$ be a localizing subcategory of $\mathcal{A}$. Then $\operatorname{Sp}(\mathcal{A}/\mathcal{X})$ is homeomorphic to the closed subset $\operatorname{Sp}\mathcal{A}\setminus\operatorname{ZSupp}\mathcal{X}$.
Proof.
We show that the image of $G:\operatorname{Sp}(\mathcal{A}/\mathcal{X})\rightarrow\operatorname{Sp}\mathcal{A}$ is $\operatorname{Sp}\mathcal{A}\setminus\operatorname{ZSupp}\mathcal{X}$. Given $I\in\operatorname{Sp}(\mathcal{A}/\mathcal{X})$, if $G(I)\in\operatorname{ZSupp}\mathcal{X}$, then there exists an object $M\in\mathcal{X}$ such that $\operatorname{Hom}_{\mathcal{A}}(M,G(I))\neq 0$. But this implies that $\operatorname{Hom}_{\mathcal{A}}(M,G(I))\cong\operatorname{Hom}_{\mathcal{A}/\mathcal{X}}(F(M),I)=0$ which is a contradiction. For every $I\in\operatorname{Sp}\mathcal{A}\setminus\operatorname{ZSupp}\mathcal{X}$, we have $I\in\mathcal{X}^{\bot}$. Since $F:\mathcal{X}^{\bot}\rightarrow\mathcal{A}/\mathcal{X}$ is a natural equivalence, $F(I)\in\operatorname{Sp}(\mathcal{A}/\mathcal{X})$ and we have $I=G(F(I))$. For every object $M^{\prime}\in\mathcal{A}/\mathcal{X}$, it is clear that we have $G(\operatorname{ZSupp}_{\mathcal{A}/\mathcal{X}}M^{\prime})=\operatorname{ZSupp}G(M^{\prime})\setminus\operatorname{ZSupp}\mathcal{X}$. On the other hand, for any $M\in\mathcal{A}$, it is clear that $G(\operatorname{ZSupp}_{\mathcal{A}/\mathcal{X}}F(M))=\operatorname{ZSupp}M\setminus\operatorname{ZSupp}\mathcal{X}$. ∎
In the rest of this section, we assume that $\mathcal{A}$ is a locally finitely generated category.
[…]
Let $M$ be an object in $\mathcal{A}$. Then
${\rm(i)}$ $\operatorname{ASupp}M=\emptyset$ if and only if $M=0$.
${\rm(ii)}$ $\operatorname{ZSupp}M=\emptyset$ if and only if $M=0$.
Proof.
(i) Since $\mathcal{A}$ is locally finitely generated, $M$ contains a finitely generated subobject $N$. Then $N$ contains a maximal subobject $K$ so that $\overline{N/K}\in\operatorname{ASupp}M$. (ii) If $M$ is nonzero, by the same notation and argument as (i), we have $E(N/K)\in\operatorname{ZSupp}N\subset\operatorname{ZSupp}M$. ∎
[…]
Let $\mathcal{X}$ be a localizing subcategory of $\mathcal{A}$ and let $M\in\mathcal{A}/\mathcal{X}$. Then $M$ is finitely generated in $\mathcal{A}/\mathcal{X}$ if and only if there exists a finitely generated object $N$ in $\mathcal{A}$ such that $M\cong N_{\mathcal{X}}$.
Proof.
Let $N$ be a finitely generated object in $\mathcal{A}$ and assume that $N_{\mathcal{X}}=\bigcup_{\lambda\in\varLambda}({M_{\lambda}})_{\mathcal{X}}$ is the direct union of subobjects in $\mathcal{A}/\mathcal{X}$. Then we have $N_{\mathcal{X}}=\bigcup_{\lambda\in\varLambda}({M_{\lambda}})_{\mathcal{X}}\subseteq(\bigcup_{\lambda\in\varLambda}{M_{i}})_{\mathcal{X}}$. On the other hand, since $N_{\mathcal{X}}=(N/t_{\mathcal{X}}(N))_{\mathcal{X}}$ and $N/t_{\mathcal{X}}(N)$ is finitely generated, we may assume that $t_{\mathcal{X}}(N)=0$. Hence
$N\cong\operatorname{Im}\eta_{N}\subseteq G(N_{\mathcal{X}})\subseteq G((\bigcup_{\lambda\in\varLambda}{M_{\lambda}})_{\mathcal{X}})=GF(\bigcup_{\lambda\in\varLambda}{M_{\lambda}})=\bigcup_{\lambda\in\varLambda}GF({M_{\lambda}}).$
The last equality holds as $GF:\mathcal{A}\rightarrow\mathcal{X}^{\bot}$ is an exact functor and it is the left adjoint of the inclusion funtor $\mathcal{X}^{\bot}\subseteq\mathcal{A}$ by [GL, Proposition 2.2]. Since $N$ is finitely generated, there exists some $\lambda\in\varLambda$ such that $N\cong\operatorname{Im}\eta_{N}\subseteq GF(M_{\lambda})$. Hence $N_{\mathcal{X}}\subseteq(M_{\lambda})_{\mathcal{X}}$; and consequently $N_{\mathcal{X}}$ is finitely generated. Let $M$ be a finitely generated object in $\mathcal{A}/\mathcal{X}$. Then there exists an object $N$ in $\mathcal{A}$ such that $M\cong N_{\mathcal{X}}$; see [GL, Proposition 2.2]. Since $\mathcal{A}$ is locally finitely generated, $N=\bigcup_{\gamma\in\varGamma}N_{\gamma}$ is the direct union of its finitely generated subobjects in $\mathcal{A}$. Since $(-)_{\mathcal{X}}$ is exact and a left adjoint functor, it preserves colimits so that $M\cong N_{\mathcal{X}}=\bigcup_{\gamma\in\varGamma}(N_{\gamma})_{\mathcal{X}}$. Hence the assumption implies that $M=(N_{\gamma})_{\mathcal{X}}$ for some $\gamma\in\varGamma.$
[…]
Let $\mathcal{X}$ be a localizing subcategory of $\mathcal{A}$. Then $\mathcal{X}$ and $\mathcal{A}/\mathcal{X}$ are locally finitely generated.
Proof.
It is clear that $\operatorname{fg-}\mathcal{X}=\operatorname{fg-}\mathcal{A}\cap\mathcal{X}$ and hence $\mathcal{X}$ is locally finitely generated. On the other hand, given an object $M$ in $\mathcal{A}/\mathcal{X}$, there exists an object $N$ in $\mathcal{A}$ such that $M=N_{\mathcal{X}}$. Since $\mathcal{A}$ is locally finitely generated, $N$ is the direct union of its finitely generated in $\mathcal{A}$. Moreover, since the localization functor $(-)_{\mathcal{X}}$ is exact and preserves colimits, it follows from section 3 that $M$ is the direct union of its finitely generated in $\mathcal{A}/\mathcal{X}$. ∎
[…]
The map $\mathcal{X}\mapsto\operatorname{ASupp}\mathcal{X}$ provides a bijective correspondence between the lattice of localizing subcategories of $\mathcal{A}$ and the lattice of open subsets of $\operatorname{ASpec}\mathcal{A}$. The inverse map is given by $\mathcal{U}\mapsto\operatorname{ASupp}^{-1}\mathcal{U}$.
Proof.
Assume that $\mathcal{X}$ is a localizing subcategory of $\mathcal{A}$ and $\mathcal{U}$ is an open subset of $\operatorname{ASpec}\mathcal{A}$. It is clear that $\operatorname{ASupp}\mathcal{X}$ is an open subset of $\operatorname{ASpec}\mathcal{A}$ and $\operatorname{ASupp}^{-1}\mathcal{U}$ is a localizing subcategory of $\mathcal{A}$. In order to prove $\operatorname{ASupp}^{-1}(\operatorname{ASupp}\mathcal{X})=\mathcal{X}$, it suffices to show that $\operatorname{ASupp}^{-1}(\operatorname{ASupp}\mathcal{X})\subset\mathcal{X}$. Given $M\in\operatorname{ASupp}^{-1}(\operatorname{ASupp}\mathcal{X})$, we have $\operatorname{ASupp}M\subset\operatorname{ASupp}\mathcal{X}$. Then $\operatorname{ASupp}M_{\mathcal{X}}=\operatorname{ASupp}M\setminus\operatorname{ASupp}\mathcal{X}=\emptyset$ so that $M_{\mathcal{X}}=0$ by section 3 and section 3. Therefore $M\in\mathcal{X}$. The fact that $\operatorname{ASupp}(\operatorname{ASupp}^{-1}\mathcal{U})=\mathcal{U}$ is straightforward. ∎
[…]
The set $\{\operatorname{ZSupp}M|\hskip 2.84544ptM\in\operatorname{fg-}\mathcal{A}\}$ forms a basis for $\operatorname{Sp}\mathcal{A}$.
Proof.
Since $\mathcal{A}$ is locally finitely generated, every $I\in\operatorname{Sp}\mathcal{A}$ contains a finitely generated subobject $M$ so that $I\in\operatorname{ZSupp}M$. If $M_{1}$ and $M_{2}$ are finitely generated objects in $\mathcal{A}$ and $I\in\operatorname{ZSupp}M_{1}\cap\operatorname{ZSupp}M_{2}$, then there exists a nonzero morphism $f_{i}:M_{i}\rightarrow I$ for $i=1,2$. Since $I$ is uniform, $\operatorname{Im}f_{1}\cap\operatorname{Im}f_{2}$ is a nonzero subobject of $I$ and so it contains a nonzero finitely generated subobject $M$. Then $I\in\operatorname{ZSupp}M\subset\operatorname{ZSupp}M_{1}\cap\operatorname{ZSupp}M_{2}$. ∎
[…]
The map $\mathcal{X}\mapsto\operatorname{ZSupp}\mathcal{X}$ provides a bijective correspondence between the lattice of localizing subcategories of $\mathcal{A}$ and the lattice of open subsets of $\operatorname{Sp}\mathcal{A}$. The inverse map is given by $\mathcal{U}\mapsto\operatorname{ZSupp}^{-1}\mathcal{U}$.
Proof.
Assume that $\mathcal{X}$ is a localizing subcategory of $\mathcal{A}$ and $\mathcal{U}$ is an open subset of $\operatorname{Sp}\mathcal{A}$. We observe that $\operatorname{ZSupp}\mathcal{X}$ is an open subset of $\operatorname{Sp}\mathcal{A}$ and $\operatorname{ASupp}^{-1}\mathcal{U}$ is a localizing subcategory of $\mathcal{A}$ by section 3. In order to prove $\operatorname{ZSupp}^{-1}(\operatorname{ZSupp}\mathcal{X})=\mathcal{X}$, it suffices to show that $\operatorname{ZSupp}^{-1}(\operatorname{ZSupp}\mathcal{X})\subset\mathcal{X}$. Given $M\in\operatorname{ZSupp}^{-1}(\operatorname{ZSupp}\mathcal{X})$, we have $\operatorname{ZSupp}M\subset\operatorname{ZSupp}\mathcal{X}$. Since $M_{\mathcal{X}}\in\mathcal{A}/\mathcal{X}$, we have $\operatorname{ZSupp}_{\mathcal{A}/\mathcal{X}}M_{\mathcal{X}}\subseteq\operatorname{Sp}(\mathcal{A}/\mathcal{X})$. Hence, it follows from section 3 that $G(\operatorname{ZSupp}_{\mathcal{A}/\mathcal{X}}M_{\mathcal{X}})=\operatorname{ZSupp}M\setminus\operatorname{ZSupp}\mathcal{X}=\emptyset$ so that $\operatorname{ZSupp}_{\mathcal{A}/\mathcal{X}}M_{\mathcal{X}}=\emptyset$. Thus $M_{\mathcal{X}}=0$ by section 3 and section 3; and consequently $M\in\mathcal{X}$. The fact that $\operatorname{ASupp}(\operatorname{ASupp}^{-1}\mathcal{U})=\mathcal{U}$ is straightforward. ∎
4. The atom spectrum of a locally coherent category
We recall that a Grothendieck category $\mathcal{A}$ is locally coherent if every object in $\mathcal{A}$ is a direct limit of coherent objects; or equivalently finitely generated subobjects of finitely presented objects are finitely presented. According to [Ro, 2] and [H] a Grothendieck category $\mathcal{A}$ is locally coherent if and only if fp-$\mathcal{A}=$coh-$\mathcal{A}$ is an abelian category. A localizing subcategory $\mathcal{X}$ of $\mathcal{A}$ is said to be of finite type provide that the corresponding right adjoint of the inclusion functor $\mathcal{X}\rightarrow\mathcal{A}$ commutes with direct limits. If $\mathcal{A}$ is a locally noetherian Grothendieck category, then fg-$\mathcal{A}=$noeth-$\mathcal{A}=$fp-$\mathcal{A}=$coh-$\mathcal{A}$ so that $\mathcal{A}$ is locally coherent. In this case, any localizing subcategory $\mathcal{X}$ of $\mathcal{A}$ is of finite type.
Throughout this section $\mathcal{A}$ is a locally coherent category. Herzog [H] associated to each $M\in\operatorname{fp-}\mathcal{A}$ a subset $\mathcal{O}(M):=\operatorname{ZSupp}M$ of $\operatorname{Sp}\mathcal{A}$ and proved that the collection $\{\mathcal{O}(M)|\hskip 2.84544ptM\in\operatorname{fp-}\mathcal{A}\}$ satisfies the axioms for a basis of open subsets defining a topology on $\operatorname{Sp}\mathcal{A}$. The resulting topological space is denoted Zg$(\mathcal{A})$, called the Ziegler spectrum of $\mathcal{A}$. In our investigation, we use the following lemma due to Krause [Kr, Lemma 1.1].
[…]

### v2

@@ 2. Categories with enough atoms
If $\sigma$ is a limit ordinal, then $\mathcal{A}_{\sigma}$ is generated by $\bigcup_{\rho<\sigma}\mathcal{A}_{\rho}$ and so we may assume that $\sigma$ is a non-limit ordinal. Let $\mathcal{C}$ be the subclass of all $\delta$-critical objects in $\mathcal{A}$ with $\delta\leq\sigma$ and we have to show that $\mathcal{A}_{\sigma}=\left\langle\mathcal{C}\right\rangle_{\rm loc}$. We prove the claim by transfinite induction on $\sigma$. The case $\sigma=0$ is clear and so we assume that $\sigma>0$. Let $\mathcal{D}$ be the class of $\sigma$-critical objects in $\mathcal{A}$. Then we have the following equalities
$F_{\sigma-1}(\mathcal{A}_{\sigma})=\left\langle F_{\sigma-1}(C)|\hskip 2.84544ptF_{\sigma-1}(C)\hskip 2.84544pt{\rm is\hskip 2.84544ptsimple\hskip 2.84544ptfor\hskip 2.84544pt}C\in\mathcal{A}\right\rangle_{\rm loc}=F_{\sigma-1}(\left\langle\mathcal{A}_{\sigma-1}\cup\mathcal{D}\right\rangle_{\rm loc})=F_{\sigma-1}(\left\langle\mathcal{C}\right\rangle_{\rm loc})$
where the first equality holds by the definition and the second equality holds by [K, Proposition 4.18] and section˜2 and the last equality holds using the induction hypothesis. It now follows from [K, Proposition 4.14] that $\mathcal{A}_{\sigma}=\left\langle\mathcal{C}\right\rangle_{\rm loc}$. ∎
The following lemma is known for modules (for example see [GW, Lemma 15.8]); but since the dimension is slightly different, we reprove it.
Lemma 2.17.
If $M$ is a nonzero object in $\mathcal{A}$ with Krull-Gabriel dimension, then $M$ has a nonzero critical subobject (and so a nonzero monoform subobject)
@@
Since ordinals satisfy the descending chain condition, we can choose a nonzero subobject $N$ of $M$ of minimal Krull-Gabriel dimension $\sigma$. Clearly $\sigma$ is non-limit ordinal and $t_{\sigma-1}(N)=0$. Since $F_{\sigma-1}(N)\in(\mathcal{A}/\mathcal{A}_{\sigma-1})_{0}$, it follows from [St, Chap VI, Proposition 2.5] that $F_{\sigma-1}(N)$ contains a simple subobject $S$. Then $N$ contains a subobject $H$ such that $F_{\sigma-1}(H)=S$ by [Po, Chap 4, Corollary 3.10]. Now, section˜2 implies that $H$ is a $\sigma$-critical. ∎
Corollary 2.18.
@@
If $M$ has Krull-Gabriel dimension, then every nonzero quotient of $M$ has Krull-Gabriel dimension as well. Thus section˜2 implies that every quotient of $M$ contains a monoform subobject so that $M$ is locally monoform. The second assertion is clear. ∎
Proposition 2.19.
The category $\mathcal{A}$ is semi-noetherian if and only if every nonzero object contains a nonzero critical subobject.
Proof.
If $\mathcal{A}$ is semi-noetherian, by section˜2, every nonzero module has a nonzero critical subobject. Conversely if $\mathcal{A}_{\tau}$ is a proper localizing subcategory of $\mathcal{A}$, then $\mathcal{A}/\mathcal{A}_{\tau}$ contains a nonzero object $Y$. Then there exists a nonzero object $X$ in $\mathcal{A}$ such that $t_{\tau}(X)=0$ and $F_{\tau}(X)=Y$ where $F_{\tau}:\mathcal{A}\rightarrow\mathcal{A}/\mathcal{A}_{\tau}$ is the canonical functor. Hence, by the assumption, $X$ contains a nonzero $\lambda+1$-critical subobject $C$, where $\lambda\geq\tau$. Consequently, by section˜2, we have $\mathcal{A}_{\tau}\neq\mathcal{A}_{\lambda+1}$ which is a contardiction. ∎
3. The atom spectrum of a locally coherent category
We recall that a Grothendieck category $\mathcal{A}$ is locally coherent if every object in $\mathcal{A}$ is a direct limit of coherent objects; or equivalently finitely generated subobjects of finitely presented objects are finitely presented. According to [Ro, 2] and [H] a Grothendieck category $\mathcal{A}$ is locally coherent if and only if fp-$\mathcal{A}=$coh-$\mathcal{A}$ is an abelian category. A localizing subcategory $\mathcal{X}$ of $\mathcal{A}$ is said to be of finite type provide that the corresponding right adjoint of the inclusion functor $\mathcal{X}\rightarrow\mathcal{A}$ commutes with direct limits. If $\mathcal{A}$ is a locally noetherian Grothendieck category, then fg-$\mathcal{A}=$noeth-$\mathcal{A}=$fp-$\mathcal{A}=$coh-$\mathcal{A}$ so that $\mathcal{A}$ is locally coherent. In this case, any localizing subcategory $\mathcal{X}$ of $\mathcal{A}$ is of finite type. Throughout this section $\mathcal{A}$ is a locally coherent category.
In our investigation, we use the following lemma due to Krause [Kr, Lemma 1.1].
Lemma 3.1.
Lemma 3.2.
Let $M$ be an object in $\mathcal{A}$. Then $\operatorname{ASupp}M=\{\alpha\in\operatorname{ASpec}\mathcal{A}|\hskip 2.84544pt\operatorname{Hom}(M,E(\alpha))\neq 0\}$.